OpenStudy (anonymous):

y=-1/√(x^2+1) then dy/dx=

OpenStudy (anonymous):

Any ideas?

OpenStudy (jennychan12):

you can rewrite it as \[y = (x^2+1)^{-1/2}\] take dy/dx

OpenStudy (anonymous):

How do you take dy/dx?

OpenStudy (jennychan12):

\[y' = -\frac{ 1 }{ 2 } (x^2+1)^{\frac{ 1 }{ 2 }}(2x)\]

OpenStudy (anonymous):

Isn't that the derivative? I thought dy/dx was implicit differentiation?

OpenStudy (anonymous):

\[dy= f'(x) \Delta x, dy= f'(x) dx\]

OpenStudy (anonymous):

\[dx= \frac{ dy }{ f'(x) }\]

OpenStudy (anonymous):

It is not implicit differentiation. If you have: \[y=f(x)\] Then the derivative can be written in several forms such as: \[\frac{dy}{dx}=f'(x)=\frac{df}{dx}\]

OpenStudy (anonymous):

Ohhhhhh... Thank you!